OpenStudy (anonymous):

What is the integral of −8x^2−8x−4dx

OpenStudy (cwrw238):

integral of ax^n = a x^(n+1) ------- n+1

OpenStudy (anonymous):

\[\int\limits_{}^{}(a+b)dx=\int\limits_{}^{}(a)dx+\int\limits_{}^{}(b)dx\] \[\int\limits_{}^{}x^{n-1}dx=\frac{x^n}{n}\]

OpenStudy (cwrw238):

oh - don't forget the arbitrary constant +C

OpenStudy (anonymous):

i plugged that in and it did not work

OpenStudy (eyust707):

lol you cannot plug the general form in as the solution.

OpenStudy (eyust707):

You must use the general form to solve for your particular case.

OpenStudy (anonymous):

i know i used the formula, but did not get the right answer. can you help me?

OpenStudy (eyust707):

Sure. What did you get?

OpenStudy (anonymous):

-8x+C

OpenStudy (eyust707):

ax^n = a x^(n+1) ------- n+1

OpenStudy (eyust707):

The first term is −8x^2

OpenStudy (eyust707):

so a= -8 n = 2

OpenStudy (eyust707):

plug in those to the formula

OpenStudy (eyust707):

Then do the next term

OpenStudy (anonymous):

i plugged them in, but its still now working